Find the average force needed to accelerates a car weighing 500 kg from rest 72 km/hr in a distance of 25 m.

Grade:12th pass

1 Answers

Mudit
7 years ago
Given mass=500kg.    initial velocity(u)=0m/s      Final velocity(v)=72km/hr=72*1000/3600=20m/s    dist.(s)=25m
acceleration(a)=?
Using v2-u2=2as
400-0=2*a*25
400/50=a       or    a=8m/s2
Force=m*a=500*8=4000N
Plz approve if correct otherwise let me know.